karen has $2000 in two banks one is earning 6% while the other is earning 3% if karen makes $84 profit by end of the year how much did she invest one each account
Answer by addingup(3677)

Let the amount in one bank, at 6%, be x.
0.06x + 0.03(2000 - x) = 84
0.06x + 60 - 0.03x = 84
0.03x = 24
x = 800
Karen has 800 earning 6% and 2000-800 = 1200 at 3%